OpenStudy (anonymous):

HELP!!! Which expression is equivalent to ((sin^2)((pi/2)-x))/sin x?

OpenStudy (anonymous):

sin(pi/2 - x) = cos(x) cos²(x)/sin(x) = cos(x) cot(x)

OpenStudy (anonymous):

Thank you oh soo very much!! @xxferrocixx